OpenStudy (anonymous):

What is the domain of the function y = sqrt3x + 3?

hartnn (hartnn):

its \(\Large \sqrt{3x+3}\) ? for the domain to be of real numbers, the quantity under the square root sign must be non-negative, so, solve \(\Large 3x+3 \ge 0\) to get your domain :)

OpenStudy (anonymous):

The domain is \[[\infty,-1]\]